# More is better: Strong constraints on the stellar properties of LEGA-C galaxies with Prospector

### v1

@@ 3.2.2 Joint posterior distributions
In Fig. 3 we show the resulting joint posterior distributions of the main physical quantities, of a handpicked quiescent galaxy with the LEGA-C ID 15. In Fig. 3, we also present the observational data (photometry and spectroscopy), along with the best-fit model, best-fit values, and the resulted SFH. A similar information is depicted in Figure 4 but for a handpicked star-forming galaxy with the LEGA-C ID 2808. Looking at the correlations between the joint posteriors we can uncover any degeneracies among the various estimated physical quantities. We determine whether two properties are correlated or not, by calculating the Spearman’s correlation coefficient (). If is very close to -1 or 1, then the two properties are strongly anti-correlated or correlated, respectively. If is very close to 0, then there is no correlation between the two properties. The Spearman’s correlation coefficient , of each joint posterior, is given in both Fig. 3 and Fig. 4.
@@
For the star-forming galaxy in Fig. 4, we find that the posterior distributions of the main physical quantities are also very well converged. Most of the joint posteriors show either no correlation or a very weak correlation. The dust attenuation slope index hints at a positive correlation with and , but still remains weak (). A correlation between the attenuation slope and optical depth is expected (Salim & Narayanan 2020; Nagaraj et al. 2022a) and has been reported in several studies in the past that use attenuation curves derived from SED fitting (e.g. Arnouts et al. 2013; Kriek & Conroy 2013; Salmon et al. 2016; Battisti et al. 2020).
@@ 3.2.3 Spectral measurements
Following the studies of Wu et al. (2018b) and Nersesian et al. (2024), we make a comparison between the observed and model spectra in terms of their Lick indices (Worthey et al. 1994). A mismatch between the observed and synthetic spectral indices, can have a strong impact on the estimates of the physical properties. We primarily focus on the comparison of the strength of two age- and metallicity-sensitive features, the HA and Fe4383. We use the Lick indices and their associated uncertainties from the updated LEGA-C catalog (van der Wel et al. 2021), measured from the emission-corrected spectra (Gallazzi et al. in prep.). The updated Lick index measurements are consistent with the DR3 values, but the uncertainties are a factor 1.3 lower (for more information see Gallazzi et al. in prep.).
@@
In Fig. 5, we show the comparison between observed and modeled Lick indices of the two aforementioned absorption features. In both panels, we see a clear division between quiescent and star-forming galaxies. Typically, star-forming galaxies showcase a strong HA absorption line (HA Å) in their spectra, while quiescent galaxies tend to be characterized by weak HA absorption (HA Å) (e.g. Kauffmann et al. 2003a, b). Inversely, quiescent galaxies tend to be metal-rich (Fe4383 Å), while star-forming galaxies tend to be metal-poor (low Fe4383). The bimodality seen in the observed spectral features is also reproduced in the distribution of modeled absorption line strengths (Straatman et al. 2018; Wu et al. 2018b; Nersesian et al. 2024).
@@
where represents the number of galaxies per Lick index, and are the model uncertainties. We find that the SEM value for HA is 0.1 Å, and for Fe4383 is 0.09 Å. These SEM values are significantly smaller than the previously reported scatter (i.e. 1.52 Å and 1.28 Å, respectively). This suggests that the differences we observe between the model and observed indices are primarily driven by systematic effects rather than random uncertainties.
@@ 3.2.4 Gain in fitting both photometry and spectroscopy
In this section, we would like to provide a detailed assessment on the benefits in fitting simultaneously photometric and spectroscopic observations of galaxies. Many studies use multi-wavelength broadband photometry to infer the physical properties of galaxies from SED fitting. Pforr et al. (2012) showed that broadband photometry, with a coverage from the restframe UV to the restframe NIR, is sufficient to resolve the stellar properties without the need of spectroscopic observations. Furthermore, many photometric surveys, both at low and high redshift, have increased their spectral coverage in the optical and NIR wavelength regime by including a great number of broadband and narrowband filters. For instance, the COSMOS survey (Scoville et al. 2007; Laigle et al. 2016; Weaver et al. 2022), or the PAU/J-PAS (Benítez et al. 2009; Abramo et al. 2012) contain more than 20 wavebands, that in principle should enable more reliable estimates of the stellar properties of galaxies. Nersesian et al. (2024) showed that, regardless of their spectral coverage, photometric SEDs cannot resolve the various spectral features that could potentially constrain the age and metallicity of stellar populations. The recent studies by Johnson et al. (2021) and Tacchella et al. (2022) provide significant evidence on the constraining of the galaxy parameters when combining spectra with photometry.
To asses the gain in fitting simultaneously photometry and spectroscopy, we performed two more SED fitting runs with Prospector in addition to our main (or fiducial) run that was presented in Section 3.1. These two new SED fitting runs include the photometry only, and spectroscopy only fits. The same physical model and prior functions were assumed in all those fits (see Table 1), but with a few differences. Specifically, we haven’t used the dust emission model in the spectroscopy-only case, while in the photometry-only case, we changed the stellar metallicity prior function. Instead of a flat prior in logarithmic space, we used a flat prior in linear space. The reason is that when a flat prior in logarithmic space is given, lower metallicity values are given a higher weight. When fitting photometry alone, a broader parameter space needs to be explored, with broader steps for the stellar metallicities. Although, a mass-metallicity prior from empirical relations from Local Universe studies (e.g. Bundy et al. 2015) could potentially improve the metallicity measurements, we argue that it could also impose a bias on the mass-metallicity relation at the redshift epoch we are studying.
@@
Now, we show how our results differ when fitting only photometry, only spectroscopy, and both photometry and spectroscopy together. Figures 7 and 8 compare the posterior distributions from the three SED fitting runs with Prospector, of the quiescent (LEGA-C ID 15) and star-forming galaxy (LEGA-C ID 2808) respectively. From the posterior distributions shown in both figures we notice that the estimates of the physical parameters between our fiducial run and the spectroscopy-only run are consistent, especially for the stellar parameters (, , sSFR). This is indicative of spectroscopy driving the properties of the stellar populations in our fitting. We also confirm that broadband SED fitting remains a reliable method for measuring stellar mass (Gallazzi et al. 2005; Conroy 2013). Focusing on the dust attenuation , we notice that the posteriors of all fitting runs return a similar value for the quiescent galaxy (low attenuation), whereas for the star-forming galaxy (high attenuation) the posterior distribution of our fiducial run lies in between the spectroscopy-only and photometry-only runs. This suggests that when estimating the properties of dusty galaxies using SED fitting, UV-MIR broadband photometry remains crucial for constraining dust properties.
@@
The measured properties from the spectroscopy-only run are almost identical to our fiducial run, without evident offsets. The spectral index is notably the only parameter that presents a significant deviation from the mean ( dex). Focusing on stellar metallicity, we observe an excellent agreement between the two fitting runs, though a few galaxies exhibit noteworthy deviations. Further inspection of these sources reveals that most of them are missing the Mg B triplet at 5176 Å, and the Fe lines at 5270 Å, 5335 Å, and 5406 Å. This suggests that to accurately constrain stellar metallicity, it is necessary to include the Mg and all Fe lines (see Gallazzi et al. in prep., for a full discussion).
When comparing our fiducial run to the photometry-only run, we find small systematic differences for some of the parameters (e.g. , sSFR), while in other cases photometry seems to constrain the parameters quite well (e.g. ). Larger discrepancies are found for the ages of the stellar populations. We find a strong systematic offset for both the light-weight ages (0.2 dex) and the mass-weighted ages (0.28 dex). Fitting only the photometry results in the estimation of older stellar ages for the galaxies, or in other words, someone would infer that galaxies have built up half of their stellar mass at earlier epochs than the spectroscopy suggests.
@@ 3.2.5 Comparing the stellar properties recovered from duplicate observations
Figure 11 shows a comparison of the three main stellar properties, namely , , and , for galaxies with duplicate observations. We segregate the observations based on the wavelength coverage, and in particular whether the Mg B absorption line is absent or not (i.e. or ). Out of the 279 galaxies with duplicate observations in our primary sample, 90 meet the wavelength coverage criterion, 52 star forming and 38 quiescent. Star-forming galaxies are depicted with blue stars, and quiescent galaxies with red points. We also calculate the absolute residuals of each parameter, the distribution of which is shown as an inset panel.
Overall, it is clear that there is an excellent agreement between the estimated properties, with the mean value of the residual distribution taking values very close to 0. This is particularly true for the quiescent galaxies, that show very little scatter. On the other hand, star-forming galaxies show more scatter, especially in their age and metallicity estimates. We also notice that the uncertainties for the observations with absent Mg and Fe absorption lines are in general larger. From this exercise we can conclude that regardless of the spectral coverage, the stellar mass and stellar age of galaxies can be systematically retrieved within 0.21 dex and 0.24 dex, respectively. Stellar metallicity is as expected more challenging, and to be able to constrain it properly a full spectral coverage is necessary (see also Gallazzi et al. in prep.).
@@ 4.1 Derived physical properties
Figure 12 shows the distributions of the main global properties of our primary galaxy sample. Each property was estimated based on the median ( percentile) of its posterior distribution. The depicted properties are the stellar mass (), the average SFR over the last 100 Myrs (SFR100Myrs), the specific SFR100Myrs (), the mass-weighted stellar age (), the light-weighted stellar age () weighted by the bolometric luminosity, the stellar metallicity (), the diffuse dust component in the band (), the dust attenuation slope index (), and the PAH mass fraction ().555We should note here that the current estimates of are solely based on the 24 m emission (or restframe m). Acquiring mid-infrared observations with JWST in the COSMOS field, can help constraining the PAH strength for galaxies at high-redshift. covering the redshift range 0.6–1 (i.e. 2 Gyr in time). We measure a median stellar mass of , with a wide range of SFRs, from passive galaxies ( SFR100Myrs/) to highly star-forming ( SFR100Myrs/). We measure a median value of SFR100Myrs/, indicating that most galaxies in our sample are actively forming new stars. Furthermore, we measure: Gyr, and . The median values of the main physical properties along with their statistical uncertainties computed as , where is the number of galaxies within the stellar mass bin, are given in Table 2 for the primary sample (2908), star-forming (1700), and quiescent (1208) galaxies. In Table 2, we also provide the median values by dividing our primary sample in four different mass bins.
@@
A closer look at the median values of both galaxy populations in Table 2, reveals that galaxies classified as quiescent, based on their colors, are characterized by very low star formation (SFR100Myrs/) and high metallicity (), compared to star-forming galaxies. On average, quiescent galaxies in our sample are 1.09 Gyr older than the star-forming ones. In general, star-forming galaxies are characterized by significantly higher dust optical depths and shallower (grayer) attenuation slopes. As mentioned in a previous section (see Section 3.2.2), a dependence between the attenuation slope and optical depth is expected (Salim & Narayanan 2020). Dust grains can either absorb or scatter the UV and optical starlight. At large optical depths (i.e. or mag) dust absorption is more dominant than scattering, resulting in a shallower (grayer) attenuation curve. As the optical depth decreases, scattering events start to dominate over absorption, resulting in a steeper attenuation curve.
@@
Going back in Fig. 12 and focusing on the distribution of the sSFR100Myrs, we notice a primary peak in the distribution at high sSFR100Myrs (), and a secondary peak at low sSFR100Myrs (), whereas the median value of the full distribution is at . Via the boundary between the two modes of the distribution, it is possible to define quiescence. Of course, there is no precise method to determine the exact boundary between quiescence and star formation in terms of sSFR. Several studies place the boundary at (e.g. Whitaker et al. 2017; Wu et al. 2018a; Leja et al. 2019c), while others place the boundary at (e.g. Gallazzi et al. 2014) or at (Brinchmann et al. 2004; Fontanot et al. 2009; Cecchi et al. 2019; Donnari et al. 2019; Paspaliaris et al. 2023). A seems to separate very well the two modes of the distribution of the sSFR100Myrs. Therefore, our galaxy sample can be alternatively divided into active and passive galaxies, depending on is above or below . According to this classification we find 1240 quiescent and 1668 star-forming galaxies. Very similar galaxy demographics were obtained through the diagram and the definition of Muzzin et al. (2013a) (see Section 2.3).
@@ 4.2 Scaling relations
The metallicity– relation also presents a weak positive trend. Comparing the trends between star-forming and quiescent galaxies, we find that quiescent galaxies show a stronger, but still weak correlation (), compared to the star-forming ones (). At fixed , quiescent galaxies are metal rich, while star-forming galaxies are metal poor. The scaling relations presented here indicate that massive galaxies with high tend to be older and have higher metallicity. In Gallazzi et al. (in prep.), we also analyze the difference between quiescent and star-forming galaxies either at fixed or , finding that quiescent galaxies are more metal-rich than star-forming ones, especially in the low-mass regime. These results are in excellent agreement with studies in the local Universe (e.g. Scott et al. 2017; Li et al. 2018), reinforcing the understanding that such galaxies have undergone significant star formation and chemical enrichment during an early epoch in their history.

### v2

@@ 3.2.2 Joint posterior distributions
In Fig. 3 we show the resulting joint posterior distributions of the main physical quantities, of a handpicked quiescent galaxy with the LEGA-C ID 15. In Fig. 3, we also present the observational data (photometry and spectroscopy), along with the best-fit model, best-fit values, and the resulted SFH. A similar information is depicted in Figure 16 but for a handpicked star-forming galaxy with the LEGA-C ID 2808. Looking at the correlations between the joint posteriors we can uncover any degeneracies among the various estimated physical quantities. We determine whether two properties are correlated or not, by calculating the Spearman’s correlation coefficient (). If is very close to -1 or 1, then the two properties are strongly anticorrelated or correlated, respectively. If is very close to 0, then there is no correlation between the two properties. The Spearman’s correlation coefficient , of each joint posterior, is given in both Fig. 3 and Fig. 16.
@@
For the star-forming galaxy in Fig. 16, we find that the posterior distributions of the main physical quantities are also very well converged. Most of the joint posteriors show either no correlation or a very weak correlation. The dust attenuation slope index hints at a positive correlation with and , but still remains weak (). A correlation between the attenuation slope and optical depth is expected (Salim & Narayanan 2020; Nagaraj et al. 2022a) and has been reported in several studies in the past that use attenuation curves derived from SED fitting (e.g., Arnouts et al. 2013; Kriek & Conroy 2013; Salmon et al. 2016; Battisti et al. 2020).
@@ 3.2.3 Spectral measurements
Following the studies of Wu et al. (2018b) and Nersesian et al. (2024), we make a comparison between the observed and model spectra in terms of their Lick indices (Worthey et al. 1994). A mismatch between the observed and synthetic spectral indices, can have a strong impact on the estimates of the physical properties. We primarily focus on the comparison of the strength of two age- and metallicity-sensitive features, the HA and Fe4383. We use the Lick indices and their associated uncertainties from the updated LEGA-C catalog (van der Wel et al. 2021), measured from the emission-corrected spectra (Gallazzi et al., in prep.). The updated Lick index measurements are consistent with the DR3 values, but the uncertainties are a factor 1.3 lower (for more information see Gallazzi et al., in prep.).
@@
In Fig. 4, we show the comparison between observed and modeled Lick indices of the two aforementioned absorption features. In both panels, we see a clear division between quiescent and star-forming galaxies. Typically, star-forming galaxies showcase a strong HA absorption line (HA Å) in their spectra, while quiescent galaxies tend to be characterized by weak HA absorption (HA Å) (e.g., Kauffmann et al. 2003a, b). Inversely, quiescent galaxies tend to be metal-rich (Fe4383 Å), while star-forming galaxies tend to be metal-poor (low Fe4383). The bimodality seen in the observed spectral features is also reproduced in the distribution of modeled absorption line strengths (Straatman et al. 2018; Wu et al. 2018b; Nersesian et al. 2024).
@@
where represents the number of galaxies per Lick index, and are the model uncertainties. We find that the SEM value for HA is 0.1 Å, and for Fe4383 is 0.09 Å. These SEM values are significantly smaller than the previously reported scatter (i.e., 1.52 Å and 1.28 Å, respectively). This suggests that the differences we observe between the model and observed indices are primarily driven by systematic effects rather than random uncertainties.
@@ 3.2.4 Gain in fitting both photometry and spectroscopy
In this section, we would like to provide a detailed assessment on the benefits in fitting simultaneously photometric and spectroscopic observations of galaxies. Many studies use multiwavelength broadband photometry to infer the physical properties of galaxies from SED fitting. Pforr et al. (2012) showed that broadband photometry, with a coverage from the restframe UV to the restframe NIR, is sufficient to resolve the stellar properties without the need of spectroscopic observations. Furthermore, many photometric surveys, both at low and high redshift, have increased their spectral coverage in the optical and NIR wavelength regime by including a great number of broadband and narrowband filters. For instance, the COSMOS survey (Scoville et al. 2007; Laigle et al. 2016; Weaver et al. 2022), or the PAU/J-PAS (Benítez et al. 2009; Abramo et al. 2012) contain more than 20 wavebands, that in principle should enable more reliable estimates of the stellar properties of galaxies. Nersesian et al. (2024) showed that, regardless of their spectral coverage, photometric SEDs cannot resolve the various spectral features that could potentially constrain the age and metallicity of stellar populations. The recent studies by Johnson et al. (2021) and Tacchella et al. (2022) provide significant evidence on the constraining of the galaxy parameters when combining spectra with photometry.
To asses the gain in fitting simultaneously photometry and spectroscopy, we performed two more SED fitting runs with Prospector in addition to our main (or fiducial) run that was presented in Section 3.1. These two new SED fitting runs include the photometry only, and spectroscopy only fits. The same physical model and prior functions were assumed in all those fits (see Table 1), but with a few differences. Specifically, we have not used the dust emission model in the spectroscopy-only case, while in the photometry-only case, we changed the stellar metallicity prior function. Instead of a flat prior in logarithmic space, we used a flat prior in linear space. The reason is that when a flat prior in logarithmic space is given, lower metallicity values are given a higher weight. When fitting photometry alone, a broader parameter space needs to be explored, with broader steps for the stellar metallicities. Although, a mass-metallicity prior from empirical relations from Local Universe studies (e.g., Bundy et al. 2015) could potentially improve the metallicity measurements, we argue that it could also impose a bias on the mass-metallicity relation at the redshift epoch we are studying.
@@
Now, we show how our results differ when fitting only photometry, only spectroscopy, and both photometry and spectroscopy together. Figures 6 and 17 compare the posterior distributions from the three SED fitting runs with Prospector, of the quiescent (LEGA-C ID 15) and star-forming galaxy (LEGA-C ID 2808) respectively. From the posterior distributions shown in both figures we notice that the estimates of the physical parameters between our fiducial run and the spectroscopy-only run are consistent, especially for the stellar parameters (, , sSFR). This is indicative of spectroscopy driving the properties of the stellar populations in our fitting. We also confirm that broadband SED fitting remains a reliable method for measuring stellar mass (Gallazzi et al. 2005; Conroy 2013). Focusing on the dust attenuation , we notice that the posteriors of all fitting runs return a similar value for the quiescent galaxy (low attenuation), whereas for the star-forming galaxy (high attenuation) the posterior distribution of our fiducial run lies in between the spectroscopy-only and photometry-only runs. This suggests that when estimating the properties of dusty galaxies using SED fitting, UV-MIR broadband photometry remains crucial for constraining dust properties.
@@
The measured properties from the spectroscopy-only run are almost identical to our fiducial run, without evident offsets. The spectral index is notably the only parameter that presents a significant deviation from the mean ( dex). Focusing on stellar metallicity, we observe an excellent agreement between the two fitting runs, though a few galaxies exhibit noteworthy deviations. Further inspection of these sources reveals that most of them are missing the Mg B triplet at 5176 Å, and the Fe lines at 5270 Å, 5335 Å, and 5406 Å. This suggests that to accurately constrain stellar metallicity, it is necessary to include the Mg and all Fe lines (see Gallazzi et al., in prep. for a full discussion).
When comparing our fiducial run to the photometry-only run, we find small systematic differences for some of the parameters (e.g., , sSFR), while in other cases photometry seems to constrain the parameters quite well (e.g., ). Larger discrepancies are found for the ages of the stellar populations. We find a strong systematic offset for both the light-weight ages (0.2 dex) and the mass-weighted ages (0.28 dex). Fitting only the photometry results in the estimation of older stellar ages for the galaxies, or in other words, someone would infer that galaxies have built up half of their stellar mass at earlier epochs than the spectroscopy suggests.
@@ 3.2.5 Comparing the stellar properties recovered from duplicate observations
Figure 9 shows a comparison of the three main stellar properties, namely , , and , for galaxies with duplicate observations. We segregate the observations based on the wavelength coverage, and in particular whether the Mg B absorption line is absent or not (i.e., or ). Out of the 279 galaxies with duplicate observations in our primary sample, 90 meet the wavelength coverage criterion, 52 star forming and 38 quiescent. Star-forming galaxies are depicted with blue stars, and quiescent galaxies with red points. We also calculate the absolute residuals of each parameter, the distribution of which is shown as an inset panel.
Overall, it is clear that there is an excellent agreement between the estimated properties, with the mean value of the residual distribution taking values very close to 0. This is particularly true for the quiescent galaxies, that show very little scatter. On the other hand, star-forming galaxies show more scatter, especially in their age and metallicity estimates. We also notice that the uncertainties for the observations with absent Mg and Fe absorption lines are in general larger. From this exercise we can conclude that regardless of the spectral coverage, the stellar mass and stellar age of galaxies can be systematically retrieved within 0.21 dex and 0.24 dex, respectively. Stellar metallicity is as expected more challenging, and to be able to constrain it properly a full spectral coverage is necessary (see also Gallazzi et al., in prep.).
@@ 4.1 Derived physical properties
Figure 10 shows the distributions of the main global properties of our primary galaxy sample. Each property was estimated based on the median ( percentile) of its posterior distribution. The depicted properties are the stellar mass (), the average SFR over the last 100 Myrs (SFR100Myrs), the specific SFR100Myrs (), the mass-weighted stellar age (), the light-weighted stellar age () weighted by the bolometric luminosity, the stellar metallicity (), the diffuse dust component in the band (), the dust attenuation slope index (), and the PAH mass fraction ().666We should note here that the current estimates of are solely based on the 24 m emission (or restframe m). Acquiring mid-infrared observations with JWST in the COSMOS field, can help constraining the PAH strength for galaxies at high-redshift. covering the redshift range 0.6–1 (i.e., 2 Gyr in time). We measure a median stellar mass of , with a wide range of SFRs, from passive galaxies ( SFR100Myrs/) to highly star-forming ( SFR100Myrs/). We measure a median value of SFR100Myrs/, indicating that most galaxies in our sample are actively forming new stars. Furthermore, we measure: Gyr, and . The median values of the main physical properties along with their statistical uncertainties computed as , where is the number of galaxies within the stellar mass bin, are given in Table 2 for the primary sample (2908), star-forming (1700), and quiescent (1208) galaxies. In Table 2, we also provide the median values by dividing our primary sample in four different mass bins.
@@
A closer look at the median values of both galaxy populations in Table 2, reveals that galaxies classified as quiescent, based on their colors, are characterized by very low star formation (SFR100Myrs/) and high metallicity (), compared to star-forming galaxies. On average, quiescent galaxies in our sample are 1.09 Gyr older than the star-forming ones. In general, star-forming galaxies are characterized by significantly higher dust optical depths and shallower (grayer) attenuation slopes. As mentioned in a previous section (see Section 3.2.2), a dependence between the attenuation slope and optical depth is expected (Salim & Narayanan 2020). Dust grains can either absorb or scatter the UV and optical starlight. At large optical depths (i.e., or mag) dust absorption is more dominant than scattering, resulting in a shallower (grayer) attenuation curve. As the optical depth decreases, scattering events start to dominate over absorption, resulting in a steeper attenuation curve.
@@
Going back in Fig. 10 and focusing on the distribution of the sSFR100Myrs, we notice a primary peak in the distribution at high sSFR100Myrs (), and a secondary peak at low sSFR100Myrs (), whereas the median value of the full distribution is at . Via the boundary between the two modes of the distribution, it is possible to define quiescence. Of course, there is no precise method to determine the exact boundary between quiescence and star formation in terms of sSFR. Several studies place the boundary at (e.g., Whitaker et al. 2017; Wu et al. 2018a; Leja et al. 2019c), while others place the boundary at (e.g., Gallazzi et al. 2014) or at (Brinchmann et al. 2004; Fontanot et al. 2009; Cecchi et al. 2019; Donnari et al. 2019; Paspaliaris et al. 2023). A seems to separate very well the two modes of the distribution of the sSFR100Myrs. Therefore, our galaxy sample can be alternatively divided into active and passive galaxies, depending on is above or below . According to this classification we find 1240 quiescent and 1668 star-forming galaxies. Very similar galaxy demographics were obtained through the diagram and the definition of Muzzin et al. (2013a) (see Section 2.3).
@@ 4.2 Scaling relations
The metallicity– relation also presents a weak positive trend. Comparing the trends between star-forming and quiescent galaxies, we find that quiescent galaxies show a stronger, but still weak correlation (), compared to the star-forming ones (). At fixed , quiescent galaxies are metal rich, while star-forming galaxies are metal poor. The scaling relations presented here indicate that massive galaxies with high tend to be older and have higher metallicity. In Gallazzi et al. (in prep.), we also analyze the difference between quiescent and star-forming galaxies either at fixed or , finding that quiescent galaxies are more metal-rich than star-forming ones, especially in the low-mass regime. These results are in excellent agreement with studies in the local Universe (e.g., Scott et al. 2017; Li et al. 2018), reinforcing the understanding that such galaxies have undergone significant star formation and chemical enrichment during an early epoch in their history.